OpenStudy (anonymous):

Can someone check my answers? Calculate the distance between the points (1,5) and (2,-3). Round your answer to the nearest tenth. Answer: -15.00 or -15? Or am I wrong all together? When calculating the midpoint of the points (3,1) and (11,7), what is the x-value of the midpoint? Answer: Not sure? When calculating the midpoint of the points (3,1) and (11,7), what is the y-value of the midpoint? Answer: 4 ?

OpenStudy (hunus):

 The distance formula is just the Pythagorean theorem. So the distance here is the hypotenuse of the triangle formed which is \[\sqrt{(2-1)^2+(-3 - (-5))^2} = \sqrt{(1)^2+(-8))^2} = \sqrt{64 + 1} = \sqrt{65}\]

OpenStudy (anonymous):

Is that all for question one or what?

OpenStudy (hunus):

The mid-point formula is \[(\frac{x_{1}+x_{2}}{2}, \frac{y_1 + y_2}{2})\] Where x1 and y1 are the coordinates of one point and x2 y2 are the coordinates of the other Notice that (x1 + x2)/2 is just the average of the two x coordinates because the average of two numbers is the number that lies between them So for the second part just plug in your x1 coordinate and x2 coordinate to find the x coordinate of the mid point